# An Adaptive Online Smoother with Closed-Form Solutions and Information-Theoretic Lag Selection for Conditional Gaussian Nonlinear Systems

### v1

@@ Appendix A Sufficient Assumptions for the Results of the CGNS Framework
$\mathbf{f}^{\mathbf{x}}(t,\mathbf{x}):=(f^{\mathbf{x}}_{1}(t,x),\ldots,f^{\mathbf{x}}_{k}(t,x))^{\mathtt{T}},\quad\mathbf{f}^{\mathbf{y}}(t,\mathbf{x}):=(f^{\mathbf{y}}_{1}(t,x),\ldots,f^{\mathbf{y}}_{l}(t,x))^{\mathtt{T}},$
$\boldsymbol{\Lambda}^{\mathbf{x}}(t,\mathbf{x}):=\left(\Lambda^{\mathbf{x}}_{ij}(t,\mathbf{x})\right)_{k\times l},\quad\boldsymbol{\Sigma}_{1}^{\mathbf{x}}(t,\mathbf{x}):=\left(\Sigma^{x,1}_{ij}(t,\mathbf{x})\right)_{k\times d},\quad\boldsymbol{\Sigma}_{2}^{\mathbf{x}}(t,\mathbf{x}):=\left(\Sigma^{x,2}_{ij}(t,\mathbf{x})\right)_{k\times r},$
$\boldsymbol{\Lambda}^{\mathbf{y}}(t,\mathbf{x}):=\left(\Lambda^{\mathbf{y}}_{ij}(t,\mathbf{x})\right)_{l\times l},\quad\boldsymbol{\Sigma}_{1}^{\mathbf{y}}(t,\mathbf{x}):=\left(\Sigma^{y,1}_{ij}(t,\mathbf{x})\right)_{l\times d},\quad\boldsymbol{\Sigma}_{2}^{\mathbf{y}}(t,\mathbf{x}):=\left(\Sigma^{y,2}_{ij}(t,\mathbf{x})\right)_{l\times r}.$
As to be able to obtain the main results which define the CGNS framework and its potency in DA through closed-form expressions for the posterior statistics, a set of sufficient regularity conditions needs to be assumed a-priori. For the following conditions, each respective pair of indices $i$ and $j$ takes all admissible values and $\mathbf{z}$ is a $k$-dimensional function in $C^{0}([0,T];\mathbb{C}^{k})$:
  1. (1)
@@
    $\int_{0}^{T}\Big{[}$ $|f^{\mathbf{y}}_{i}(t,\mathbf{z})|+|\Lambda^{\mathbf{y}}_{ij}(t,\mathbf{z})|+\big{|}f^{\mathbf{x}}_{i}(t,\mathbf{z})\big{|}^{2}+\big{|}\Lambda^{\mathbf{x}}_{ij}(t,\mathbf{z})\big{|}^{2}$
    $+\big{|}\Sigma^{x,1}_{ij}(t,\mathbf{z})\big{|}^{2}+\big{|}\Sigma^{x,2}_{ij}(t,\mathbf{z})\big{|}^{2}+\big{|}\Sigma^{y,1}_{ij}(t,\mathbf{z})\big{|}^{2}+\big{|}\Sigma^{y,2}_{ij}(t,\mathbf{z})\big{|}^{2}\Big{]}\mathrm{d}t<+\infty.$
@@
  2. (2)
    If $g(t,\mathbf{z})$ denotes any element of the multiplicative factors in the drift dynamics, $\boldsymbol{\Lambda}^{\mathbf{x}}$ and $\boldsymbol{\Lambda}^{\mathbf{y}}$, or of the noise feedback matrices, $\boldsymbol{\Sigma}_{m}^{\mathbf{x}}(t,\mathbf{z})$ and $\boldsymbol{\Sigma}_{m}^{\mathbf{y}}(t,\mathbf{z})$ for $m=1,2$, and $K(s)$ is a nondecreasing right-continuous function taking values in $[0,1]$, then there $\exists L_{2},L_{3},L_{4},L_{5}>0$ such that for any $\mathbf{z}$ and $\mathbf{w}$ $k$-dimensional functions in $C^{0}([0,T];\mathbb{C}^{k})$ we have
@@
    $|g(t,\mathbf{z})|^{2}\leq L_{4}\int_{0}^{t}(1+\left\|\mathbf{z}(s)\right\|_{2}^{2})\mathrm{d}K(s)+L_{5}(1+\left\|\mathbf{z}(t)\right\|_{2}^{2}),\ \forall t\in[0,T].$
@@
  3. (3)
    The sum of the Gramians (with respect to rows) of the noise coefficient matrices in the observable process are uniformly nonsingular, i.e., the elements of the inverse of
@@
  4. (4)
    $\int_{0}^{T}\mathbb{E}\left[|\Lambda^{\mathbf{x}}_{ij}(t,\mathbf{z}(t))y_{j}(t)|\right]\mathrm{d}t<+\infty$.
  5. (5)
    $\mathbb{E}\left[|y_{j}(t)|\right]<+\infty,\ t\in[0,T]$.
  6. (6)
    For ${\mu_{\text{f}}}_{j}(t):=\mathbb{E}\left[y_{j}(t)\big{|}\mathcal{F}_{t}^{\mathbf{x}}\right]$, where $t\in[0,T]$ and $j=1,\ldots,l$, we assume that
    $\mathbb{P}\left(\int_{0}^{T}\left|\Lambda^{\mathbf{x}}_{ij}(t,\mathbf{z}(t))\mu_{\text{f},j}(t)\right|^{2}\mathrm{d}t<+\infty\right)=1.$
Accompanying the conditional Gaussianity provided by Theorem 2.1, we further assume the following additional conditions (which concur with assumptions (1)(6) from before):
@@
    $|\Lambda^{\mathbf{x}}_{ij}(t,\mathbf{z})|,|\Lambda^{\mathbf{y}}_{ij}(t,\mathbf{z})|\leq L_{1}$ for some (uniform) $L_{1}>0$, $\forall t\in[0,T]$.
@@
    $\int_{0}^{T}\mathbb{E}\left[|\boldsymbol{\Lambda}_{ij}^{\mathbf{x}}(t,\mathbf{x})|^{4}+|\mathbf{f}_{i}^{\mathbf{y}}(t,\mathbf{x})|^{4}+|\Sigma^{y,1}_{ij}(t,\mathbf{x})|^{4}+|\Sigma^{y,2}_{ij}(t,\mathbf{x})|^{4}\right]\mathrm{d}t<+\infty$.
@@
Finally, as to be able to derive discrete-time smoother equations which are consistent with their continuous-time counterparts in the limit as their associated temporal partition’s mesh size converges to zero, with this recovery being first order-accurate under an explicit Euler-Maruyama numerical scheme, we assume that:
@@
    $\mathbb{E}\left[\|\mathbf{x}(0)\|_{2}^{2}+\|\mathbf{y}(0)\|_{2}^{2}\right]<+\infty$. This also implies $\mathbb{P}$-a.s. finiteness of the initial conditions by Markov’s inequality.
  • (11)
    For $\mathbf{z}\in C^{0}([0,T];\mathbb{C}^{k})$, $\mathbf{u}\in C^{0}([0,T];\mathbb{C}^{l})$, and $\mathbf{f}:=\begin{pmatrix}\mathbf{0}_{k\times k}&\boldsymbol{\Lambda}^{\mathbf{x}}\\ \mathbf{0}_{l\times k}&\boldsymbol{\Lambda}^{\mathbf{y}}\end{pmatrix}\begin{pmatrix}\mathbf{x}\\ \mathbf{y}\end{pmatrix}+\begin{pmatrix}\mathbf{f}^{\mathbf{x}}\\ \mathbf{f}^{\mathbf{y}}\end{pmatrix}$, $\boldsymbol{\Sigma}:=\begin{pmatrix}\boldsymbol{\Sigma}_{1}^{\mathbf{x}}&\boldsymbol{\Sigma}_{2}^{\mathbf{x}}\\ \boldsymbol{\Sigma}_{2}^{\mathbf{y}}&\boldsymbol{\Sigma}_{2}^{\mathbf{y}}\end{pmatrix}$, there exist $L_{6},L_{7}>0$ such that
    $\left\|\mathbf{f}(t,\mathbf{z},\mathbf{u})-\mathbf{f}(s,\mathbf{z},\mathbf{u})\right\|_{2}+\left\|\boldsymbol{\Sigma}(t,\mathbf{z})-\boldsymbol{\Sigma}(s,\mathbf{z})\right\|_{2}\leq L_{6}\big{(}1+(\left\|\mathbf{z}\right\|^{2}_{2}+\left\|\mathbf{u}\right\|^{2}_{2})^{1/2}\big{)}|t-s|^{1/2},\ \forall t,s\in[0,T].$
[…]
      We implicitly assume that condition (10) from Andreou & Chen [68] is satisfied for any stochastic discrete-time approximation which recovers the strong solution to the CGNS at vanishing mesh size, and specifically for the numerical solution provided by the Euler-Maruyama scheme. In other words, we assume that the initial distributions are not being severely misidentified.
  • (12)
    Adopting a uniform temporal discretization of $[0,T]$, with $\Delta t=T/J$ being its norm or mesh and $t_{j}=j\Delta t$ for $j=0,1,\ldots,J$ being its nodes, where $J\in\mathbb{N}$, we assume that
@@
The optimal discrete smoother equations derived under these additional sufficient assumptions are essential in establishing a corresponding optimal online discrete smoother which runs forward in time.
@@ Appendix B Proof of the Optimal Online Forward-In-Time Discrete Smoother
$\boldsymbol{\mu}_{\text{{s}}}^{j,n}$ $=\boldsymbol{\mu}_{\text{{f}}}^{j}+\mathbf{E}^{j}\left(\boldsymbol{\mu}_{\text{{s}}}^{j+1,n}-(\mathbf{I}_{l\times l}+\boldsymbol{\Lambda}^{\mathbf{y},j}\Delta t)\boldsymbol{\mu}_{\text{{f}}}^{j}-\mathbf{f}^{\mathbf{y},j}\Delta t\right)+\mathbf{F}^{j}\left(\mathbf{x}^{j+1}-\mathbf{x}^{j}-(\boldsymbol{\Lambda}^{\mathbf{x},j}\boldsymbol{\mu}_{\text{{f}}}^{j}+\mathbf{f}^{\mathbf{x},j})\Delta t\right)=\mathbf{E}^{j}\boldsymbol{\mu}_{\text{{s}}}^{j+1,n}+\mathbf{b}^{j},$ (B.1)
@@
$\mathbf{b}^{j}$ $:=\boldsymbol{\mu}_{\text{f}}^{j}-\mathbf{E}^{j}\left((\mathbf{I}_{l\times l}+\boldsymbol{\Lambda}^{\mathbf{y},j}\Delta t)\boldsymbol{\mu}_{\text{{f}}}^{j}+\mathbf{f}^{\mathbf{y},j}\Delta t\right)+\mathbf{F}^{j}\left(\mathbf{x}^{j+1}-\mathbf{x}^{j}-(\boldsymbol{\Lambda}^{\mathbf{x},j}\boldsymbol{\mu}_{\text{{f}}}^{j}+\mathbf{f}^{\mathbf{x},j})\Delta t\right),$
@@
which immediately shows that (3.14) and (3.15) hold by plugging-in $j=n-1$ into these, since the smoother and filter estimates are the same at the end point, i.e., $\boldsymbol{\mu}_{\text{s}}^{n,n}=\boldsymbol{\mu}_{\text{f}}^{n}$ and $\mathbf{R}_{\text{s}}^{n,n}=\mathbf{R}_{\text{f}}^{n}$. As such, we just need to prove (3.11) and (3.12).
@@
The curved arrow pointing to the right above the product symbol indicates the order with which we expand said product since we are dealing with matrices, thus making the product, in general, noncommutative. Also, note that (B.3) is able to encompass both the second and third relations of (3.13), but not the first one due to its irregular form. Nonetheless, it could be extraneously incorporated in (B.3) by noting that if $j=n$, and so the product on the right-hand side is “empty”, then $\mathbf{D}^{n,n-1}$ is trivially defined as the identity matrix $\mathbf{I}_{l\times l}$.
@@
$\mathbf{D}^{j,n-1}=\mathbf{D}^{j,n-2}\mathbf{E}^{n-1}$ =$=\begin{cases}\mathbf{E}^{n-1},&\text{ for }j=n-1\\ \mathbf{E}^{n-2}\mathbf{E}^{n-1},&\text{ for }j=n-2\\ \mathbf{D}^{j,n-3}\mathbf{E}^{n-2}\mathbf{E}^{n-1},&\text{ for }j<n-2=\begin{cases}\mathbf{E}^{n-3}\mathbf{E}^{n-2}\mathbf{E}^{n-1},&\text{ for }j=n-3\\ \mathbf{D}^{j,n-4}\mathbf{E}^{n-3}\mathbf{E}^{n-2}\mathbf{E}^{n-1},&\text{ for}j<n-3\end{cases}\end{cases}$
@@
We proceed now by proving that for $n\in\mathbb{N}$ and $10\leq j\leq n-$ 1, the following equations are valid:
@@
$=\cdots\ \left(\text{Likewise where at the }k\text{-th step we have }\boldsymbol{\mu}_{\text{s}}^{j+k,n}=\begin{cases}\mathbf{E}^{j+k}\boldsymbol{\mu}_{\text{s}}^{j+k+1,n}+\mathbf{b}^{j+k},&\text{ for }j+k<n\\ \boldsymbol{\mu}_{\text{f}}^{n,n},&\text{ for }j+k=n\end{cases}\right)$
@@
$=\mathbf{E}^{j}\mathbf{E}^{j+1}\cdots\mathbf{E}^{n-1}\boldsymbol{\mu}_{\text{f}}^{n}+\mathbf{b}^{j}+\left(\mathbf{E}^{j}\mathbf{b}^{j+1}+\mathbf{E}^{j}\mathbf{E}^{j+1}\mathbf{b}^{j+2}+\cdots+\left(\overset{{\curvearrowright}}{\prod^{n-2}_{i=j}}\mathbf{E}^{i}\right)\mathbf{b}^{n-1}\right).$
@@
$=\cdots\ \left(\text{Likewise where at the }k\text{-th step we have }\boldsymbol{\mu}_{\text{s}}^{j+k,n}=\begin{cases}\mathbf{E}^{j+k}\boldsymbol{\mu}_{\text{s}}^{j+k+1,n}+\mathbf{b}^{j+k},&\text{ for }j+k<n\\ \boldsymbol{\mu}_{\text{f}}^{n,n},&\text{ for }j+k=n\end{cases}\right)$
@@
$\hskip 56.9055pt+\left(\mathbf{E}^{j}\mathbf{P}^{j+1}_{j+2}(\mathbf{E}^{j})^{\dagger}+\mathbf{E}^{j}\mathbf{E}^{j+1}\mathbf{P}^{j+2}_{j+3}(\mathbf{E}^{j+1})^{\dagger}(\mathbf{E}^{j})^{\dagger}+\cdots+\left(\overset{{\curvearrowright}}{\prod^{n-2}_{i=j}}\mathbf{E}^{i}\right)\mathbf{P}^{n-1}_{n}\left(\overset{{\curvearrowleft}}{\prod^{n-2}_{i=j}}(\mathbf{E}^{i})^{\dagger}\right)\right).$
@@
$(\mathbf{D}^{j,r-1})^{\dagger}=\left(\overset{{\curvearrowright}}{\prod^{r-1}_{i=j}}\mathbf{E}^{i}\right)^{\dagger}=\overset{{\curvearrowleft}}{\prod^{r-1}_{i=j}}(\mathbf{E}^{i})^{\dagger},$
@@
$=\mathbf{R}_{\text{s}}^{j,n-1}+\mathbf{D}^{j,n-2}\left(\mathbf{R}_{\text{s}}^{n-1,n}-\mathbf{R}_{\text{f}}^{n-1}\right),$

### v2

@@ Appendix A Sufficient Assumptions for the Results of the CGNS Framework
$\mathbf{f}^{\mathbf{x}}(t,\mathbf{x}):=\big(f^{\mathbf{x}}_{1}(t,\mathbf{x}),\ldots,f^{\mathbf{x}}_{k}(t,\mathbf{x})\big)^{\mathtt{T}},\quad\mathbf{f}^{\mathbf{y}}(t,\mathbf{x}):=\big(f^{\mathbf{y}}_{1}(t,\mathbf{x}),\ldots,f^{\mathbf{y}}_{l}(t,\mathbf{x})\big)^{\mathtt{T}},$
$\boldsymbol{\Lambda}^{\mathbf{x}}(t,\mathbf{x}):=\big(\Lambda^{\mathbf{x}}_{ij}(t,\mathbf{x})\big)_{k\times l},\quad\boldsymbol{\Sigma}_{1}^{\mathbf{x}}(t,\mathbf{x}):=\big(\Sigma^{\mathbf{x},1}_{ij}(t,\mathbf{x})\big)_{k\times d},\quad\boldsymbol{\Sigma}_{2}^{\mathbf{x}}(t,\mathbf{x}):=\big(\Sigma^{\mathbf{x},2}_{ij}(t,\mathbf{x})\big)_{k\times r},$
$\boldsymbol{\Lambda}^{\mathbf{y}}(t,\mathbf{x}):=\big(\Lambda^{\mathbf{y}}_{ij}(t,\mathbf{x})\big)_{l\times l},\quad\boldsymbol{\Sigma}_{1}^{\mathbf{y}}(t,\mathbf{x}):=\big(\Sigma^{\mathbf{y},1}_{ij}(t,\mathbf{x})\big)_{l\times d},\quad\boldsymbol{\Sigma}_{2}^{\mathbf{y}}(t,\mathbf{x}):=\big(\Sigma^{\mathbf{y},2}_{ij}(t,\mathbf{x})\big)_{l\times r}.$
To be able to obtain the main results which define the CGNS framework and its potency in DA through closed-form expressions for the posterior statistics, a set of sufficient regularity conditions needs to be assumed a-priori. In what follows, each respective pair of indices $i$ and $j$ takes all admissible values and $\mathbf{z}$ is a $k$-dimensional function in $C^{0}\big([0,T];\mathbb{C}^{k}\big)$:
  • (1)
@@
    $\int_{0}^{T}\Big[$ $\big|f^{\mathbf{y}}_{i}(t,\mathbf{z})\big|+\big|\Lambda^{\mathbf{y}}_{ij}(t,\mathbf{z})\big|+\big|f^{\mathbf{x}}_{i}(t,\mathbf{z})\big|^{2}+\big|\Lambda^{\mathbf{x}}_{ij}(t,\mathbf{z})\big|^{2}$
    $+\big|\Sigma^{\mathbf{x},1}_{ij}(t,\mathbf{z})\big|^{2}+\big|\Sigma^{\mathbf{x},2}_{ij}(t,\mathbf{z})\big|^{2}+\big|\Sigma^{\mathbf{y},1}_{ij}(t,\mathbf{z})\big|^{2}+\big|\Sigma^{\mathbf{y},2}_{ij}(t,\mathbf{z})\big|^{2}\Big]\mathrm{d}t<+\infty.$
@@
  • (2)
    $\big|\Lambda^{\mathbf{x}}_{ij}(t,\mathbf{z})\big|,\big|\Lambda^{\mathbf{y}}_{ij}(t,\mathbf{z})\big|\leq L_{1}$ for some (uniform) $L_{1}>0$, $\forall t\in[0,T]$.
  • (3)
    If $g(t,\mathbf{z})$ denotes any element of the multiplicative factors in the drift dynamics, $\boldsymbol{\Lambda}^{\mathbf{x}}$ and $\boldsymbol{\Lambda}^{\mathbf{y}}$, or of the noise feedback matrices, $\boldsymbol{\Sigma}_{m}^{\mathbf{x}}(t,\mathbf{z})$ and $\boldsymbol{\Sigma}_{m}^{\mathbf{y}}(t,\mathbf{z})$ for $m=1,2$, and $K(s)$ is a nondecreasing right-continuous function taking values in $[0,1]$, then there $\exists L_{2},L_{3},L_{4},L_{5}>0$ such that for any $\mathbf{z}$ and $\mathbf{w}$ $k$-dimensional functions in $C^{0}\big([0,T];\mathbb{C}^{k}\big)$ we have:
@@
    $|g(t,\mathbf{z})|^{2}\leq L_{4}\int_{0}^{t}\big(1+\left\|\mathbf{z}(s)\right\|_{2}^{2}\big)\mathrm{d}K(s)+L_{5}\big(1+\left\|\mathbf{z}(t)\right\|_{2}^{2}\big),\ \forall t\in[0,T].$
@@
  • (4)
    $\mathbb{E}\big[\|\mathbf{x}(0)\|_{2}^{2}+\|\mathbf{y}(0)\|_{2}^{2}\big]<+\infty$.
  • (5)
    The sum of the Gramians (with respect to the rows) of the noise coefficient matrices in the observable process are uniformly nonsingular, i.e., the elements of the inverse of
@@
    $\int_{0}^{T}\mathbb{E}\left[\big|\Lambda^{\mathbf{x}}_{ij}(t,\mathbf{z}(t))y_{j}(t)\big|\right]\mathrm{d}t<+\infty$.
@@
    $\mathbb{E}\Big[\big|y_{j}(t)\big|\Big]<+\infty,\ t\in[0,T]$.
  • (8)
    For $\mu_{\text{f},j}(t):=\mathbb{E}\big[y_{j}(t)\big|\mathcal{F}_{t}^{\mathbf{x}}\big]$, where $t\in[0,T]$ and $j=1,\ldots,l$, we assume that:
    $\mathbb{P}\left(\int_{0}^{T}\big|\Lambda^{\mathbf{x}}_{ij}(t,\mathbf{z}(t))\mu_{\text{f},j}(t)\big|^{2}\mathrm{d}t<+\infty\right)=1,$
    for all $i=1,\ldots,k$.
  • (9)
    For $\mathbf{z}\in C^{0}\big([0,T];\mathbb{C}^{k}\big)$, $\mathbf{u}\in C^{0}\big([0,T];\mathbb{C}^{l}\big)$, and $\mathbf{f}:=\begin{pmatrix}\mathbf{0}_{k\times k}&\boldsymbol{\Lambda}^{\mathbf{x}}\\ \mathbf{0}_{l\times k}&\boldsymbol{\Lambda}^{\mathbf{y}}\end{pmatrix}\begin{pmatrix}\mathbf{x}\\ \mathbf{y}\end{pmatrix}+\begin{pmatrix}\mathbf{f}^{\mathbf{x}}\\ \mathbf{f}^{\mathbf{y}}\end{pmatrix}$, $\boldsymbol{\Sigma}:=\begin{pmatrix}\boldsymbol{\Sigma}_{1}^{\mathbf{x}}&\boldsymbol{\Sigma}_{2}^{\mathbf{x}}\\ \boldsymbol{\Sigma}_{1}^{\mathbf{y}}&\boldsymbol{\Sigma}_{2}^{\mathbf{y}}\end{pmatrix}$, there exists $L_{6}>0$ such that
    $\left\|\mathbf{f}(t,\mathbf{z},\mathbf{u})-\mathbf{f}(s,\mathbf{z},\mathbf{u})\right\|_{2}+\left\|\boldsymbol{\Sigma}(t,\mathbf{z})-\boldsymbol{\Sigma}(s,\mathbf{z})\right\|_{2}\leq L_{6}\left(1+\big(\left\|\mathbf{z}\right\|^{2}_{2}+\left\|\mathbf{u}\right\|^{2}_{2}\big)^{1/2}\right)|t-s|^{1/2},\ \forall t,s\in[0,T].$
  • (10)
    Let $\big(\mathbf{x}_{\Delta t}(t),\mathbf{y}_{\Delta t}(t)\big)$ be the stochastic discrete-time approximation of $(\mathbf{x},\mathbf{y})$, after a continuous-time extension, that is provided by the numerical integration Euler–Maruyama scheme when applied on the CGNS in (2.1)–(2.2) (see (3.1)–(3.2)), over a temporal uniform partition of $[0,T]$ with $\Delta t$ being the uniform subinterval length. Then, as in Andreou & Chen [55], we assume that there exists $L_{7}>0$ such that:
    $\mathbb{E}\left[\big\|\mathbf{x}(0)-\mathbf{x}_{\Delta t}(0)\big\|^{2}_{2}+\big\|\mathbf{y}(0)-\mathbf{y}_{\Delta t}(0)\big\|^{2}_{2}\right]\leq L_{7}\Delta t.$
    In other words, we assume that the exact initial distributions are not being severely misidentified by the discrete-time ones (at least in the mean).
  • (11)
    $\int_{0}^{T}\mathbb{E}\left[\big|\boldsymbol{\Lambda}_{ij}^{\mathbf{x}}(t,\mathbf{x})\big|^{4}+\big|\mathbf{f}_{i}^{\mathbf{y}}(t,\mathbf{x})\big|^{4}+\big|\Sigma^{\mathbf{y},1}_{ij}(t,\mathbf{x})\big|^{4}+\big|\Sigma^{\mathbf{y},2}_{ij}(t,\mathbf{x})\big|^{4}\right]\mathrm{d}t<+\infty$.
@@
    As with (10), by adopting a uniform temporal discretization of $[0,T]$, with $\Delta t=T/J$ being its norm and $t_{j}=j\Delta t$ for $j=0,1,\ldots,J$ being its nodes, where $J\in\mathbb{N}$, we assume that
@@
For more details about the regularity conditions in (1)(13), including possible relaxations, see Andreou & Chen [55]. Most importantly, in the context of this work, these assumptions are sufficient for establishing the CGNS optimal online forward-in-time discrete smoother from Theorem 3.1.
@@ Appendix B Proof of the Optimal Online Forward-In-Time Discrete Smoother
$\boldsymbol{\mu}_{\text{{s}}}^{j,n}$ $=\boldsymbol{\mu}_{\text{{f}}}^{j}+\mathbf{E}^{j}\big(\boldsymbol{\mu}_{\text{{s}}}^{j+1,n}-(\mathbf{I}_{l\times l}+\boldsymbol{\Lambda}^{\mathbf{y},j}\Delta t)\boldsymbol{\mu}_{\text{{f}}}^{j}-\mathbf{f}^{\mathbf{y},j}\Delta t\big)+\mathbf{F}^{j}\big(\mathbf{x}^{j+1}-\mathbf{x}^{j}-(\boldsymbol{\Lambda}^{\mathbf{x},j}\boldsymbol{\mu}_{\text{{f}}}^{j}+\mathbf{f}^{\mathbf{x},j})\Delta t\big)=\mathbf{E}^{j}\boldsymbol{\mu}_{\text{{s}}}^{j+1,n}+\mathbf{b}^{j},$ (B.1)
@@
$\mathbf{b}^{j}$ $:=\boldsymbol{\mu}_{\text{f}}^{j}-\mathbf{E}^{j}\big((\mathbf{I}_{l\times l}+\boldsymbol{\Lambda}^{\mathbf{y},j}\Delta t)\boldsymbol{\mu}_{\text{{f}}}^{j}+\mathbf{f}^{\mathbf{y},j}\Delta t\big)+\mathbf{F}^{j}\big(\mathbf{x}^{j+1}-\mathbf{x}^{j}-(\boldsymbol{\Lambda}^{\mathbf{x},j}\boldsymbol{\mu}_{\text{{f}}}^{j}+\mathbf{f}^{\mathbf{x},j})\Delta t\big),$
@@
which immediately shows that (3.13) and (3.14) hold by plugging in j=n1j=n-1 into these, since the smoother and filter estimates are the same at the end point, i.e., $\boldsymbol{\mu}_{\text{s}}^{n,n}=\boldsymbol{\mu}_{\text{f}}^{n}$ and $\mathbf{R}_{\text{s}}^{n,n}=\mathbf{R}_{\text{f}}^{n}$. As such, we just need to prove (3.10) and (3.11).
@@
The curved arrow pointing to the right above the product symbol in (B.3) indicates the order or direction with which we expand said product, since we are dealing with matrices, thus making the product, in general, noncommutative. Also, note that (B.3) is able to encompass both the second and third relations of (3.12), but not the first one due to its irregular form. Nonetheless, it could be extraneously incorporated in (B.3) by noting that if $j=n$, and so the product on the right-hand side is “empty”, then $\mathbf{D}^{n,n-1}$ is trivially defined as the identity matrix $\mathbf{I}_{l\times l}$.
@@
[…]
@@
We proceed now by proving that for $n\in\mathbb{N}$ and $0\leq j\leq n-1$, the following equations are valid:
@@
 $\hskip 5.69046pt\vdots\hskip 22.76228pt\text{ (Inductively...)}$
$=\mathbf{E}^{j}\mathbf{E}^{j+1}\cdots\mathbf{E}^{n-1}\boldsymbol{\mu}_{\text{f}}^{n}+\mathbf{b}^{j}+\Bigg(\mathbf{E}^{j}\mathbf{b}^{j+1}+\mathbf{E}^{j}\mathbf{E}^{j+1}\mathbf{b}^{j+2}+\cdots+\Bigg(\overset{{\curvearrowright}}{\prod^{n-2}_{i=j}}\mathbf{E}^{i}\Bigg)\mathbf{b}^{n-1}\Bigg).$
@@
$\hskip 56.9055pt+\Bigg(\mathbf{E}^{j}\mathbf{P}^{j+1}_{j+2}(\mathbf{E}^{j})^{\dagger}+\mathbf{E}^{j}\mathbf{E}^{j+1}\mathbf{P}^{j+2}_{j+3}(\mathbf{E}^{j+1})^{\dagger}(\mathbf{E}^{j})^{\dagger}+\cdots+\Bigg(\overset{{\curvearrowright}}{\prod^{n-2}_{i=j}}\mathbf{E}^{i}\Bigg)\mathbf{P}^{n-1}_{n}\Bigg(\overset{{\curvearrowleft}}{\prod^{n-2}_{i=j}}(\mathbf{E}^{i})^{\dagger}\Bigg)\Bigg).$
@@
$(\mathbf{D}^{j,r-1})^{\dagger}=\Bigg(\overset{{\curvearrowright}}{\prod^{r-1}_{i=j}}\mathbf{E}^{i}\Bigg)^{\dagger}=\overset{{\curvearrowleft}}{\prod^{r-1}_{i=j}}(\mathbf{E}^{i})^{\dagger},$
@@
$=\mathbf{R}_{\text{s}}^{j,n-1}+\mathbf{D}^{j,n-2}\left(\mathbf{R}_{\text{s}}^{n-1,n}-\mathbf{R}_{\text{f}}^{n-1}\right)(\mathbf{D}^{j,n-2})^{\dagger},$